OpenStudy (anonymous):

\[\frac{ e ^{x ^{2}} - 1 }{ x ^{2} }\] derived

OpenStudy (anonymous):

@ganeshie8 @UnkleRhaukus @Hero

OpenStudy (anonymous):

u have to Find Derivative right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Use Quotient Rule..

OpenStudy (anonymous):

\[e ^{2x} / 2x\] Is correct?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

then

OpenStudy (anonymous):

First Write the Quotient Rule..

OpenStudy (anonymous):

\[x ^{2} e ^{2x} - e ^{x ^{2}-1} * 2x / (x ^{2})^{2}\]

OpenStudy (shubhamsrg):

do you,,by any chance, have this around in your question ?  its just my intuition ..

OpenStudy (anonymous):

yes

OpenStudy (shubhamsrg):

lol.. :D and by derive , you mean application of L hopital rule right ?

OpenStudy (anonymous):

if x=1 how to solve this derived

OpenStudy (anonymous):

f = e^x^2 - 1 f' = (e^x^2)(2x) = 2xe^x^2 g = x^2 g' = 2x (f' g - g' f)/ g^2 = ((2) (e^x^2) (x^2) - (2) (e^x^2) + 2) / x^3